# H-Fac: Memory-Efficient Optimization with Factorized Hamiltonian Descent

### v1

@@ Abstract
In this study, we introduce a novel adaptive optimizer, H-Fac, which incorporates a factorized approach to momentum and scaling parameters. Our algorithm demonstrates competitive performance on both ResNets and Vision Transformers, while achieving sublinear memory costs through the use of rank-1 parameterizations for moment estimators. We develop our algorithms based on principles derived from Hamiltonian dynamics, providing robust theoretical underpinnings. These optimization algorithms are designed to be both straightforward and adaptable, facilitating easy implementation in diverse settings.
@@ 4.1 Adafactor optimizer as Hamiltonian
     $X_{t}=X_{t-1}-\eta_{t}\left(\mathrm{clip}\left(M_{t}/\sqrt{\hat{V}_{t}}\right)+\lambda X_{t}\right)$
@@ 4.2 Factorized first-order momentum estimation
3) When the momentum coefficient is deactivated ($\beta_{1}=0$), the update terms inside $\nabla\phi(.)$ and $\nabla\psi(.)$ degenerate to $\nabla f(X_{t})$. In this scenario, the algorithm simplifies to conventional gradient descent:

### v2

@@ Abstract
In this study, we introduce a novel adaptive optimizer, H-Fac, which incorporates a factorized approach to momentum and scaling parameters. Our algorithm demonstrates competitive performances on both ResNets and Vision Transformers, while achieving sublinear memory costs through the use of rank-1 parameterizations for moment estimators. We develop our algorithms based on principles derived from Hamiltonian dynamics, providing robust theoretical underpinnings. These optimization algorithms are designed to be both straightforward and adaptable, facilitating easy implementation in diverse settings.
@@ 4.1 Adafactor optimizer as Hamiltonian
     $X_{t}=X_{t-1}-\eta_{t}\left(\mathrm{clip}\left(M_{t}/\sqrt{\hat{V}_{t}}\right)+\lambda X_{t-1}\right)$
@@ 4.2 Factorized first-order momentum estimation
Algorithm 2 signFSGD for matrix parameter, with factored first-order moments.
  Inputs: moment coefficients $\beta=0.9$, and regularization constant $\lambda$
  Initialization: weight parameters $X_{0}\in\mathbb{R}^{m\times n}$, initial moment factors $u_{0},v_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     $u_{t}=\beta u_{t-1}+(1-\beta)G_{t}1_{n}/n$
     $v_{t}=\beta v_{t-1}+(1-\beta)G_{t}^{\top}1_{m}/m$
     $\hat{u}_{t}=u_{t}-G_{t}1_{n}/n$
     $\hat{v}_{t}=v_{t}-G_{t}^{\top}1_{m}/m$
     $X_{t}=X_{t-1}-\eta_{t}\left(\mathrm{sign}\left(\beta\hat{u}_{t}1_{n}^{\top}+G_{t}\right)+\mathrm{sign}\left(\beta 1_{m}\hat{v}_{t}^{\top}+G_{t}\right)+\lambda X_{t-1}\right)$
  end for
@@
3) When the momentum coefficient is deactivated ($\beta_{1}=0$), the update terms inside $\nabla\phi(.)$ and $\nabla\psi(.)$ degenerate to $\nabla f(X_{t})$. In this scenario, the algorithm simplifies to standard gradient descent:
@@ Appendix B Factorized Lion optimizer
Algorithm 4 Lionfactor for matrix parameter, with factored first-order moments.
  Inputs: double-moment coefficients $\beta_{1}=0.9,\beta_{2}=0.99$, and regularization constant $\lambda$
  Initialization: weight parameters $X_{0}\in\mathbb{R}^{m\times n}$, initial moment factors $u_{0},v_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     update model parrameters
     $\hat{u}_{t}=\beta_{1}u_{t-1}+(1-\beta_{1})G_{t}1_{n}/n-G_{t}1_{n}/n$
     $\hat{v}_{t}=\beta_{1}v_{t-1}+(1-\beta_{1})G_{t}^{\top}1_{m}/m-G_{t}^{\top}1_{m}/m$
     $X_{t}=X_{t-1}-\eta_{t}\left(\mathrm{sign}\left(\hat{u}_{t}1_{n}^{\top}+G_{t}\right)+\mathrm{sign}\left(1_{m}\hat{v}_{t}^{\top}+G_{t}\right)+\lambda X_{t-1}\right)$
     update exponential moment averages
     $u_{t}=\beta_{2}u_{t-1}+(1-\beta_{2})G_{t}1_{n}/n$
     $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})G_{t}^{\top}1_{m}/m$
  end for